# Omni-Video: Democratizing Unified Video Understanding and Generation

### v2

@@ Abstract
Notable breakthroughs in unified understanding and generation modeling have led to remarkable advancements in image understanding, reasoning, production and editing, yet current foundational models predominantly focus on processing images, creating a gap in the development of unified models for video understanding and generation. This report presents Omni-Video, an efficient and effective unified framework for video understanding, generation, as well as instruction-based editing. Our key insight is to teach existing multimodal large language models (MLLMs) to produce continuous visual clues that are used as the input of diffusion decoders, which produce high-quality videos conditioned on these visual clues. To fully unlock the potential of our system for unified video modeling, we integrate several technical improvements: 1) a lightweight architectural design that respectively attaches a vision head on the top of MLLMs and a adapter before the input of diffusion decoders, the former produce visual tokens for the latter, which adapts these visual tokens to the conditional space of diffusion decoders; and 2) an efficient multi-stage training scheme that facilitates a fast connection between MLLMs and diffusion decoders with limited data and computational resources. We empirically demonstrate that our model exhibits satisfactory generalization abilities across video generation, editing and understanding tasks. Our code will be publicly available at: https://github.com/SAIS-FUXI/Omni-Video. $\dagger$ Equal Contribution. * Corresponding Authors.
@@ 5.2 Experiment Results
Performance on GenEval benchmark
@@
Table 2: Performance Comparison on text-to-image benchmark GenEval.
Method Params Single Obj. Two Obj. Count. Colors Pos. Color Attri. Overall
SDv1.5 rombach2022high 0.9B 0.97 0.38 0.35 0.76 0.04 0.06 0.43
SDv2.1 rombach2022high 0.9B 0.98 0.51 0.44 0.85 0.07 0.17 0.50
SDXL sdxl 2.6B 0.98 0.74 0.39 0.85 0.15 0.23 0.55
PixArt-$\alpha$ chen2023pixartalpha 0.6B 0.98 0.50 0.44 0.80 0.08 0.07 0.48
DALL-E 2 dalle2 6.5B 0.94 0.66 0.49 0.77 0.10 0.19 0.52
DALL-E 3 dalle3 - 0.96 0.87 0.47 0.83 0.43 0.45 0.67
Chameleon team2024chameleon 34B - - - - - - 0.39
LWM liu2024world 7B 0.93 0.41 0.46 0.79 0.09 0.15 0.47
@@
Show-o xie2024show 1.3B 0.95 0.52 0.49 0.82 0.11 0.28 0.53
JanusFlow ma2025janusflow 1.3B 0.97 0.59 0.45 0.83 0.53 0.42 0.63
Ours 11B 0.93 0.52 0.51 0.67 0.10 0.20 0.49

### v3

@@ Abstract
Notable breakthroughs in unified understanding and generation modeling have led to remarkable advancements in image understanding, reasoning, production and editing, yet current foundational models predominantly focus on processing images, creating a gap in the development of unified models for video understanding and generation. This report presents Omni-Video, an efficient and effective unified framework for video understanding, generation, as well as instruction-based editing. Our key insight is to teach existing multimodal large language models (MLLMs) to produce continuous visual clues that are used as the input of diffusion decoders, which produce high-quality videos conditioned on these visual clues. To fully unlock the potential of our system for unified video modeling, we integrate several technical improvements: 1) a lightweight architectural design that respectively attaches a vision head on the top of MLLMs and a adapter before the input of diffusion decoders, the former produce visual tokens for the latter, which adapts these visual tokens to the conditional space of diffusion decoders; and 2) an efficient multi-stage training scheme that facilitates a fast connection between MLLMs and diffusion decoders with limited data and computational resources. We empirically demonstrate that our model exhibits satisfactory generalization abilities across video generation, editing and understanding tasks. Our code is publicly available at: https://github.com/SAIS-FUXI/Omni-Video. $\dagger$ Equal Contribution. * Corresponding Authors.
@@ 5.2 Experiment Results
T2I performance on GenEval benchmark
@@
T2V performance on VBench-Long benchmark.
We evaluated our model’s long-range video generation performance on the VBench-Long huang2024vbench benchmark, which comprises 16 fine-grained metrics. As shown in Tab. 3, our method achieves state-of-the-art results in Subject Consistency (98.39%), Temporal Flickering (99.87%), Object Class recognition (93.54%), and Temporal Style consistency (25.81%). These results highlight the model’s strong ability to preserve temporal coherence, maintain subject identity, and capture semantic details over extended sequences. Compared with Wan2.1-T2V-1.3B, which reports the highest quality score (85.23%), our model delivers competitive overall performance while surpassing it in 10 out of 16 categories, demonstrating the robustness and generalization ability of our unified video generation framework.
Table 2: T2I performance comparison on GenEval.
Method Single Obj Two Obj Count Color Pos Color Attri Overall
SDv1.5 rombach2022high 0.97 0.38 0.35 0.76 0.04 0.06 0.43
SDv2.1 rombach2022high 0.98 0.51 0.44 0.85 0.07 0.17 0.50
SDXL sdxl 0.98 0.74 0.39 0.85 0.15 0.23 0.55
PixArt-$\alpha$ chen2023pixartalpha 0.98 0.50 0.44 0.80 0.08 0.07 0.48
DALL-E 2 dalle2 0.94 0.66 0.49 0.77 0.10 0.19 0.52
DALL-E 3 dalle3 0.96 0.87 0.47 0.83 0.43 0.45 0.67
Chameleon team2024chameleon - - - - - - 0.39
LWM liu2024world 0.93 0.41 0.46 0.79 0.09 0.15 0.47
@@
Show-o xie2024show 0.95 0.52 0.49 0.82 0.11 0.28 0.53
JanusFlow ma2025janusflow 0.97 0.59 0.45 0.83 0.53 0.42 0.63
Ours 0.99 0.89 0.84 0.87 0.35 0.56 0.75
Table 3: T2V performance on VBench-Long.
Overall Scores Technical Quality Aesthetic Quality
Model Name Total Score Quality Semantic Subject Background Temporal Motion Dynamic Aesthetic Imaging
EasyAnimateV5.1 fu2024easyanimate 83.42 85.03 77.01 98.00 97.41 99.19 98.02 57.15 69.48 68.61
MiniMax-Video-01 minimax2024video 83.41 84.85 77.65 97.51 97.05 99.10 99.22 64.91 63.03 67.17
Kling 1.6 kuaishou2025kling 83.40 85.00 76.99 97.40 96.84 98.64 99.13 62.22 64.81 69.70
Wan2.1-T2V-1.3B wan2025wan 83.31 85.23 75.65 97.56 97.93 99.55 98.52 65.19 65.46 67.01
HunyuanVideo kong2024hunyuanvideo 83.24 85.09 75.82 97.37 97.76 99.44 98.99 70.83 60.36 67.56
Gen-3 runway2024gen3 82.32 84.11 75.17 97.10 96.62 98.61 99.23 60.14 63.34 66.82
Vchitect-2.0 (VEnhancer) fan2025vchitect 82.24 83.54 77.06 96.83 96.66 98.57 98.98 63.89 60.41 65.35
CogVideoX1.5-5B yuan2024cogvideox 82.17 82.78 79.76 96.87 97.35 98.88 98.31 50.93 62.79 65.02
Ours 83.00 84.27 77.92 98.39 97.68 99.87 99.10 56.67 62.48 64.56
Semantic Fidelity
Model Name Object Multi-Obj Action Color Spatial Scene Appearance Temporal Overall
EasyAnimateV5.1 fu2024easyanimate 89.57 66.85 95.60 77.86 76.11 54.31 23.06 24.61 26.47
MiniMax-Video-01 minimax2024video 87.83 76.04 92.40 90.36 75.50 50.68 20.06 25.63 27.10
Kling 1.6 kuaishou2025kling 93.34 63.99 96.20 81.26 79.08 55.57 20.75 24.51 26.04
Wan2.1-T2V-1.3B wan2025wan 88.81 74.83 94.00 89.20 73.04 41.96 21.81 23.13 25.50
HunyuanVideo kong2024hunyuanvideo 86.10 68.55 94.40 91.60 68.68 53.88 19.80 23.89 26.44
Gen-3 runway2024gen3 87.81 53.64 96.40 80.90 65.09 54.57 24.31 24.71 26.69
Vchitect-2.0 (VEnhancer) fan2025vchitect 86.61 68.84 97.20 87.04 57.55 56.57 23.73 25.01 27.57
CogVideoX1.5-5B yuan2024cogvideox 87.47 69.65 97.20 87.55 80.25 52.91 24.89 25.19 27.30
Ours 93.54 71.06 93.60 88.89 73.15 44.33 23.45 25.81 26.99